THE ROLES OF ATOMIC OXYGEN AND NITRIC OXIDE IN LOW TEMPERATURE PLASMAS

\underline{SHERRIE S. BOWMAN}, DAVID BURNETTE, IGOR V. ADAMOVICH, and WALTER R. LEM-PERT, M.A. Chaszeyka Non-Equilibrium Thermodynamics Laboratory, The Ohio State University, Columbus, Ohio 43210..

Nitric oxide, NO, and oxygen, O, concentrations have been experimentally measured using one-photon and two-photon laser induced fluorescence, respectively, as a function of time after a nanosecond pulsed plasma discharge. The relative behavior of these two species is fundamentally different than that predicted using the extended Zeldovich Mechanism. A plasma chemistry kinetic model sensitivity analysis has been conducted to determine the dominate reactions involved. The spectra, concentrations, and kinetic modeling predictions will be discussed.